OpenStudy (anonymous):

five added to 3 times a number is equal to 8 less than 5 times the number

OpenStudy (anonymous):

are you in flvs??

OpenStudy (anonymous):

what is that

OpenStudy (anonymous):

5+3xa=5-a

OpenStudy (anonymous):

florida virtual school

OpenStudy (anonymous):

?

OpenStudy (anonymous):

no idaho

OpenStudy (anonymous):

iva

OpenStudy (anonymous):

so its 5+3xa=5-a you need to some how figure out a, do you have any more info?

OpenStudy (anonymous):

five added to 3 times a number is equal to 8 less than 5 times the number 5+3x=5x-8 Solve for x

OpenStudy (anonymous):

the closest one there is 4x + 3 = 8x- 5